What is the mass of 2.40 ×1022 molecules of NaOH (Molar mass = 40.0 g/mol)?
1
Expert's answer
2019-09-23T03:47:50-0400

n(NaOH) = N(NaOH) / NA = 2.40 × 1022 / 6.02 × 1023 mol-1 = 0.0399 mol.

m(NaOH) = n(NaOH) × M(NaOH) = 0.0399 mol × 40 g/mol = 1.596 g ≈ 1.60 g.